OpenStudy (anonymous):

A Farmer Leans a 12 ft Ladder Against A Barn. The Base of the Ladder Is 3ft From The Barn.l To The Nearest Foot, How High on the Barn Does The Ladder Reach??? A.) 11ft B.) 12ft. C.) 13ft. D.) 14 ft If Someone Can Answer This, Id Really Appreciate it!

OpenStudy (anonymous):

For this you just need to use the Pythagorean Theorem. \[a^2+b^2=c^2\]\[(3)^2+b^2=(12)^2\]\[(9)+b^2=(144)\]\[b^2=135\]\[b \approx11 ft.\]

OpenStudy (anonymous):

Sorry rounded incorrectly the first time. The answer is: \[b \approx12\]

OpenStudy (anonymous):

\[3\text{ Tan}\left[\text{ArcCos }\left[\frac{3}{12}\right]\right]=3 \sqrt{15}=11.619 \text{ feet}\]B.)